OpenStudy (anonymous):

write the decimal as a fraction or a mixed number in the simplest form 3.4 (the 4 has a line on the top) need someone to help me do this not only give me the answer. thanx

OpenStudy (anonymous):

like this \[3.\overline{4}\]?

OpenStudy (anonymous):

i.e. \(3.44444...\)

OpenStudy (anonymous):

i know you want a method, and i will explain one, but the easiest thing here is to know that \[\overline{.1}=.1111...=\frac{1}{9}\] and \[\overline{.2}=.2222...=\frac{2}{9}\] and \[\overline{.3}=.3333....=\frac{3}{9}=\frac{1}{3}\] and so on

OpenStudy (anonymous):

so that \(\overline{.4}=.4444...=\frac{4}{9}\) and therefore \[3\overline{.4}=3.444...=3\tfrac{4}{9}\]

OpenStudy (anonymous):

i cnt see anytthig